OpenStudy (anonymous):

(27a^6)^2/3 in simplest form please

OpenStudy (ajprincess):

\(\Large27=3^3\) \(\Large(27a^6)^{\frac{2}{3}}\) \(=\Large{27^{\frac{2}{3}}a^{6*{\frac{2}{3}}}}\) \(=\Large{3^{3*{\frac{2}{3}}}a^{6*{\frac{2}{3}}}}\) Can u find the simplest form @rhodes?

OpenStudy (anonymous):

i dont get it

OpenStudy (ajprincess):

\(27=3*3*3=3^3\) Getting this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is the answer?? ___a^_ ??

OpenStudy (ajprincess):

good:) Using this general rule \(\Large(a^b)^c=a^{b*c}\) Nw \(\Large(27a^6)^{\frac{2}{3}}\) \(=\Large27^{\frac{2}{3}}a^{6*{\frac{2}{3}}}\) Getting this?

OpenStudy (anonymous):

yes but what is the answer ___a^- ??

OpenStudy (ajprincess):

Actually I am sorry I am supposed to give u direct answers.

OpenStudy (anonymous):

ok carry on i get up to the last bit

OpenStudy (ajprincess):

U knw that \(27=3^3\) Plugging that in \(\Large27^{\frac{2}{3}}a^{6*\frac{2}{3}}\) u get \(\Large3^{3*\frac{2}{3}}a^{6*\frac{2}{3}}\) Getting this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@rhodes , simplify the last equation that ajprincess wrote...

OpenStudy (nubeer):

i don't think you are oblige to give direct answers @ajprincess , you were teaching the way to get the answer so you don't have to be sorry... @rhodes this is a learning site not an answering site.. if you want answer then this is not the place .. you should go and read code of conduct if u are about to ask question here..

OpenStudy (anonymous):

scarydoor u were a lot of help thanks :)

OpenStudy (anonymous):

uhh... I think @ajprincess might have been a bit more deserving of thanks...

OpenStudy (anonymous):

naa